OpenStudy (anonymous):

Help with Algebra 2B? Rationalize the denominator and simplify (wait for the equation in comments):

OpenStudy (anonymous):

\[\frac{ 3\sqrt{2}-2\sqrt{3} }{ 3\sqrt{2}+2\sqrt{3} }\]

OpenStudy (anonymous):

@Hero

hero (hero):

Multiply top and bottom by the conjugate.

OpenStudy (anonymous):

So @Hero it would be \[\frac{ (3\sqrt{2}-2\sqrt{3})(3\sqrt{2}-2\sqrt{3}) }{ (3\sqrt{2}+2\sqrt{3})(3\sqrt{2}-2\sqrt{3}) }\] which would simplify to \[\frac{ 3\sqrt{2}^{2}+2\sqrt{3}^{2} }{ 3\sqrt{2}^{2}-2\sqrt{3}^{2} }\] am I right so far?

hero (hero):

It's more like: \[\frac{ (3\sqrt{2}-2\sqrt{3})(3\sqrt{2}-2\sqrt{3}) }{ (3\sqrt{2}+2\sqrt{3})(3\sqrt{2}-2\sqrt{3}) } = \frac{(3\sqrt{2}-2\sqrt{3})^2}{(3\sqrt{2})^2 - (2\sqrt{3})^2}\]

OpenStudy (anonymous):

Oh ok thanks that helps a lot I think I can do it from here.